OpenStudy (anonymous):

There are two factors of -36 such that one factor is 11 less than half of the other factor. Choose all the pairs of these factors: -2 and 18 -6 and 6 3 and -12 4 and -9

OpenStudy (anonymous):

3 and -12

OpenStudy (johnweldon1993):

Can set up a system of equations... xy = -36 x = (y/2) - 11 So we substitute that into the first equation \[\large (\frac{y}{2} - 11)y = -36\] \[\large \frac{y^2}{2} - 11y = -36\] \[\large y^2 - 22y = -72\] \[\large y^2 - 22y + 121 = 49\] \[\large (y - 11)^2 = 49\] \[\large y - 11 = \sqrt{49}\] This breaks into 2 equations \[\large y - 11 = 7\] and \[\large y - 11 = -7\] Solving we get \[\large y = 18\] \[\large y = 4\] if those are 2 possible 'y' values...we can us them both to solve for 2 different 'x' values too xy = -36 x(18) = -36 x = -2 and xy = -36 x(4) = -36 x = -9 So we have 2 pairs... -2 and 18 and -9 and 4